OpenStudy (alyssajobug):

What is the solution to: √2x+13 -5=x?

OpenStudy (isaiah.feynman):

Can you use the equation editor to write the question?

OpenStudy (alyssajobug):

\[\sqrt{2x}+13-5=x\]

OpenStudy (isaiah.feynman):

What is 13-5?

OpenStudy (alyssajobug):

8

OpenStudy (alyssajobug):

\[\sqrt{2x}+8=x\]

OpenStudy (isaiah.feynman):

Subtract 8 from both sides.

OpenStudy (alyssajobug):

\[\sqrt{2x}=x -8\]

OpenStudy (alyssajobug):

I got this far, the only thing really confusing me a ton is what to do with the radical

OpenStudy (isaiah.feynman):

Relax buddy, we will destroy it!

OpenStudy (anonymous):

lol isaiah

OpenStudy (isaiah.feynman):

Square both sides to remove the radical!

OpenStudy (alyssajobug):

\[4x =x ^{2}-64?\]

OpenStudy (isaiah.feynman):

Not quite.

OpenStudy (alyssajobug):

\[2x =\left( x-8 \right)^{2}\]

OpenStudy (alyssajobug):